The blades on a fan have a frequency of 15 Hz.

a. How many cycles do the blades complete in five minutes?

b. What is the period for the fan blades?

c. How long would it take for the fan blades to complete 100 revolutions?


1
Expert's answer
2021-03-29T08:58:59-0400

(a) We can find the number of cycles as follows:


"N=5\ min\cdot\dfrac{60\ s}{1\ min}\cdot\dfrac{15\ cycles}{1\ s}=4500\ cycles."

(b) We can find the period for the fan blades as follows:


"T=\dfrac{1}{f}=\dfrac{1}{15\ Hz}=0.067\ s."

(c) We can find the time that the fan blades takes to complete 100 revolutions as follows:


"t=100\ rev\cdot\dfrac{1\ s}{15\ rev}=6.67\ s."